OpenStudy (anonymous):

Finding the length of a parametric curve. x=5sin^3t, y=5cos^3t, 0≤t≤pi I've gotten the problem to this point: 15 Integral (sqrt(cos^2t*sin^2)dt I'm not putting the bounds for convenience. I put this in the calculator and get 15, but I don't know how that happens. I know that's right because I also have the answer key, so there's that. But wouldn't what's in the integral simplify to sin t * cos t? Because it doesn't in the calculator, and that gives a different answer if you integrate that in the calculator. But it makes sense. Help?

OpenStudy (anonymous):

Basically, my question is what is \[\sqrt{\sin^2(x)\cos^2(x)}\] equal to, if it's not equal to sin(x)cos(x) Help would be the best.